### v1

# Training-Free Identity Preservation in Stylized Image Generation Using Diffusion Models

Abstract
While diffusion models have demonstrated remarkable generative capabilities, existing style transfer techniques often struggle to maintain identity while achieving high-quality stylization. This limitation is particularly acute for images where faces are small or exhibit significant camera-to-face distances, frequently leading to inadequate identity preservation. To address this, we introduce a novel, training-free framework for identity-preserved stylized image synthesis using diffusion models. Key contributions include: (1) the ”Mosaic Restored Content Image” technique, significantly enhancing identity retention, especially in complex scenes; and (2) a training-free content consistency loss that enhances the preservation of fine-grained content details by directing more attention to the original image during stylization. Our experiments reveal that the proposed approach substantially surpasses the baseline model in concurrently maintaining high stylistic fidelity and robust identity integrity, particularly under conditions of small facial regions or significant camera-to-face distances, all without necessitating model retraining or fine-tuning.
Figure 1: Overview of our identity-preserving pipeline. First, YOLOv8 [35] detects faces in the content image, and Real-ESRGAN [39] enhances them. The enhanced faces are then combined with the original background to form a mosaic image. This mosaic image undergoes style transfer, producing stylized faces and a stylized background within the mosaic. Finally, the stylized faces are extracted from this processed mosaic and reinserted into the final output to maintain identity details.
1 Introduction
In recent years, image style transfer has become a transformative technique in computer vision, leveraging deep learning and generative models to modify an image’s visual style while preserving its intrinsic content structure [19]. The core of style transfer lies in the disentangling of two distinct aspects of an image: its content and its style [7]. Content information refers to the structural elements of an image such as the spatial arrangement of objects and the identities of individuals, while style information comprises aesthetic attributes, including texture, color distribution, and edge rendering, that define its overall visual appearance [27]. In a typical style transfer task, a content image and a style reference image are provided with the goal of extracting stylistic features from the reference image and applying them to the content image, thereby altering its visual appearance while preserving its underlying structure [11]. Style transfer finds applications across multiple industries, including digital art and design for creating unique artworks, marketing and advertising for customized visual content, and entertainment for novel visual effects [1, 2].
1.1 Challenges in Diffusion-Based Style Transfer
While diffusion models such as Stable Diffusion XL [32] have made significant advancements in image generation quality, they are challenged by two primary issues. The first challenge is the Number of Inference Steps. The iterative nature of diffusion models requires multiple inference steps to achieve high fidelity in the generated image [34]. As the number of steps increases, the output quality improves; however, this also results in higher computational demands and increased processing time. The second challenge is Identity Preservation. Recent advancements, such as inversion processes that prepare initial latent representations for the UNet [29, 10], have shown promise in reducing inference steps, particularly in scenarios where we aim to perform image editing tasks on the input image and accurately reconstruct it. However, these low-step approaches often struggle to preserve critical content details, particularly the identity of individuals in the image. Our observations reveal that while facial features are maintained for subjects captured close to the camera, those that are farther away tend to lose their distinctive characteristics. To address this issue, especially for faces that are farther from the camera or in images with multiple faces, we have developed a novel technique that operates in a training-free manner, ensuring robust identity preservation. This technique significantly improves the consistency of facial features and the overall retention of identity. Previous works [37, 38], including the implementation of IP adapters [46] that inject identity information into the model via cross-attention mechanisms, have mitigated this problem to some extent. However, these methods often fail in scenarios that involve multiple faces or in which subjects are positioned at a distance. In addition to improving identity preservation, another focus of this research is the refinement of noise during the denoising steps of Stable Diffusion, which enhances the model’s ability to retain detailed content and identity features in the generated image.
In this work, we employ the Instance Style Plus pipeline [38], based on Stable Diffusion, to stylize content images using style reference images. In Chapter 2, we present a concise review of the evolution of models and techniques for image style transfer. Chapter 3 describes our proposed pipeline in detail, and Chapter 4 provides a comparative analysis of our approach against the baseline model.
2 Related Work
The field of artistic style transfer has undergone significant transformations due to the rapid advancement of generative AI techniques. A notable example is the transfer of image styles [12, 23, 25, 26, 30, 41, 49] in recent years, where the appearance of an image is altered by incorporating the artistic features of one or more reference images. The aim is to merge the style elements of the source images with the content of the target image to create stylized output. Based on advances in generative AI models, style transfer has evolved into three primary domains: Neural Style Transfer (NST) [12], Generative Adversarial Networks (GANs) [13], and Diffusion Models [16]. In the following, we will explore the evolution of these approaches, highlighting their strengths and limitations, and ultimately focusing on the domain most relevant to our work: diffusion-based style transfer.
2.1 Neural Style Transfer (NST)
The traditional Neural Style Transfer (NST) approach to image style transfer relies on non-parametric texture transformation methods [15, 5] but often struggles with complex images due to its reliance on fundamental, rather than abstract, feature extraction. Gatys et al.[12]represent the style and content of an image as two distinct parts by combining convolutional neural networks (CNN) with texture generation methods. In this method, the high-level convolutional layers of the CNN capture the image’s content, while the low-level layers represent the style. However, this approach faces challenges such as slow convergence, and insufficient stylistic output. Johnson et al.[20] improved the efficiency of style transfer by employing feedforward networks with perceptual loss functions and a pretrained VGG model. However, similar to [12], their approach faced difficulties in achieving realistic results due to the limitations of the Gram matrix in style extraction.
2.2 GANs
GANs have become fundamental to artistic stylization due to their ability to capture high-level data patterns. They are widely applied in various tasks, including painting [8], cartooning [9], and font stylization [43]. More GAN-based approaches [13, 18, 21, 31, 50]integrate paired [18] and unpaired [21, 31, 50] datasets to improve model flexibility and generalization. The early methods relied on conditional GAN architectures, such as Pix2Pix [18]. These methods are supervised, meaning they need paired examples of input and output images. In contrast, CycleGAN [50] proposes an unsupervised image-to-image translation model that learns bidirectional mappings between two domains using adversarial loss and cycle consistency. This approach enables effective style transfer without the need for paired training data. DualStyleGAN [45] proposed a dual-path model that allows control of both structural and artistic styles. Furthermore, DCT-Net [27] introduced a few-shot portrait stylization model that enhances style transfer by first calibrating the content distribution and then applying localized translation.
2.3 Diffusion Models
Recent advancements in diffusion models have revolutionized image stylization, enabling sophisticated and adaptable style transfer techniques [3]. Some methods leverage textual inversion, with InST [48] using textual inversion to encode styles as text embeddings to offer more flexible style representation. Stylediffusion [42] introduced a CLIP-based loss for better style-content disentanglement during diffusion model training. Furthermore, several approaches utilize text as a conditioning mechanism for style control or content generation [44, 6]. Training-free methods, such as DiffStyle [45] modify skip connections within h-space [22] to balance style and content preservation. However, when used with Stable Diffusion [36], it can sometimes change textures and the overall layout of the image in ways that aren’t intended. A major challenge within style transfer maintaining the integrity of the original content; for instance, a portrait intended to resemble a painting must still depict the same person. StyleID [4] tries to solve this by adjusting how the image is processed and using some clever tricks to preserve the original content with query preservation and initial latent AdaIN.
inversion-based models like StyleAlign [14]which uses shared self-attention and Adaptive Instance Normalization (AdaIN)[17] to get an accurate style transfer but these methods may struggle with style-content leakage. More recent works, like IP-Adapter [46] and Style-Adapter [40] offer greater control and flexibility in style transfer by using a specialized cross-attention mechanism that processes text and image features in separate attention layers. Despite their advantages, encoder-based methods like IP-Adapter face challenges in differentiating style from content. Therefore, ControlNet [47] is employed to effectively regulate and maintain the content of the generated image, ensuring that the content remains intact while applying the desired stylistic transformation. InstantStyle [37] enhances IP-Adapter for image stylization by incorporating the CLIP embedding of the style images into targeted blocks within SDXL. This method effectively separates content and style, enabling a more precise and adaptable style transfer process. InstantStyle-Plus [38] introduces a novel diffusion-based style transfer method, employing a tile-based ControlNet, semantic adapter, and style extractor to improve content preservation and style fidelity. This study demonstrates that by utilizing specific components, the performance of diffusion models in style transfer can be significantly enhanced.
3 Methodology
Figure 2: Overview of our proposed style transfer pipeline. Our method builds upon the InstantStyle-Plus pipeline [38] by incorporating a Mosaic Restored Content Image to preserve identity in images. To further enhance identity preservation and detail fidelity, we introduce Content Consistency Loss, which refines the denoising process.
In the Introduction section, we identified two key challenges in the style transfer process. The first challenge is the need for a high number of inference steps to achieve high-quality image synthesis with diffusion models. This results in increased computational demands and longer processing times, which may limit their practicality in industrial settings. While this issue can be partially addressed by employing the pipeline introduced in [38] along with inversion algorithms [10], it leads to a secondary problem: the failure to preserve the identity and details of individuals in the scene when the number of inference steps is reduced. In our research, we investigated the impact of facial size and the number of subjects on identity preservation, particularly focusing on scenarios where faces are small or images contain multiple individuals. Our experiments reveal that the identity preservation performance of the base pipeline is dependent on the size and quantity of faces within the image. Specifically, when faces are sufficiently large, individual identity details are well maintained; however, when the facial regions are small or the image contains multiple faces, a significant loss of distinctive identity features is observed. To address this issue, we propose a solution that incorporates face detection and image enhancement techniques. Specifically, we used yolo [35] face detection model to identify the faces of individuals in the scene. Subsequently, we employ Real-ESRGAN image enhancement model [39] to improve the clarity and quality of the detected faces. In [33], it was observed that using image restoration techniques can significantly enhance image segmentation by making objects more clearly defined. Therefore, to increase the overall image quality, we incorporate this technique into our pipeline. The next step involves creating a mosaic image that contains the enhanced face images and the original background. This mosaic image is used as an input to the style transfer pipeline, ensuring that facial identities are preserved. In the final output, the stylized face images are extracted from the stylized mosaic image and placed back into their original positions in the scene, keeping the integrity of the content image (see Figure 1).
This approach, which utilizes straightforward computational methods, offers a robust solution to the identity preservation issue; specifically, it significantly improves the retention of facial identities in stylized output.
3.1 Improving the Denoising Process for Better Content Preservation
In image stylization using Stable Diffusion, a key challenge is balancing the transfer of style with the preservation of content details, such as identity and structure. In text-to-image models like SDXL, which are trained on billions of images across diverse styles, the text prompt plays a pivotal role in determining the output’s stylistic domain. Consequently, when employing targeted style keywords to generate images in a specific style, it becomes necessary to increase the model’s attention to the text prompt. However, this heightened focus on style can inadvertently compromise content fidelity, leading to the loss of critical details, particularly the identity of objects or individuals. While adding control networks like Canny, tile, or depth maps can improve detail retention, they also increase the model’s size, slow down processing, and consume more memory. Despite this, even with additional control networks, the model still struggles to preserve details effectively.
To address this problem, we propose a refined denoising process for latents in VAE space that preserves the identity and structure of the content while still allowing effective style transfer. The key idea is to refine the predicted noise at each timestep and ensure that the latents remain close to their original content representation. This technique involves denoising the initial latent by refining the predicted noise at each step, allowing us to maintain a balance between content and style. Our method focuses on calculating the loss between the stylized image (created with the predicted noise) and the original content image, both of which exist in VAE space. We then compute the gradient of this loss with respect to the predicted noise and use this gradient to iteratively refine the noise, bringing the latents closer to their initial latent content.
The process begins with the generation of an initial stylized image in the VAE space, where we compute the L2 distance between the stylized image and the original content image. This loss guides the model to preserve the details of the content while applying the style. The gradient of this loss is used to adjust the predicted noise, which is defined in multiple steps.
3.2 Mathematical Formulation
Given the latent representation $z_{t}$ in the timestep $t$, initialized from a content image via an inversion algorithm [10], we estimate the denoised content representation using the following.
$\hat{x}_{0}=\frac{z_{t}-\sqrt{1-\alpha_{t}}\cdot\hat{\epsilon}_{t}}{\sqrt{\alpha_{t}}}$ (1)
where $\hat{x}_{0}$ represents the estimated denoised image in vae space, $\hat{\epsilon}_{t}$ is the predicted noise at timestep $t$, and $\alpha_{t}$ is the cumulative product of noise schedule parameters. To ensure that the stylized output remains close to the original content, we define the content consistency loss as:
$\mathcal{L}_{\text{content}}=\|\hat{x}_{0}-x_{c}\|_{2}^{2}$ (2)
where $x_{c}$ is the original content latent representation in vae space. To refine the noise prediction, we compute the gradient of the content consistency loss with respect to the predicted noise:
$\frac{\partial\mathcal{L}_{\text{content}}}{\partial\hat{\epsilon}_{t}}=2(\hat{x}_{0}-x_{c})\cdot\left(-\frac{\sqrt{1-\alpha_{t}}}{\sqrt{\alpha_{t}}}\right)$ (3)
Using a hyperparameter $\lambda_{c}$ to control the refinement strength, the refined predicted noise is computed as:
$\hat{\epsilon}_{t}^{\text{refined}}=\hat{\epsilon}_{t}-\lambda_{c}\cdot\frac{\partial\mathcal{L}_{\text{content}}}{\partial\hat{\epsilon}_{t}}$ (4)
where $\lambda_{c}$ is a tunable hyperparameter that determines the strength of the content preservation constraint. Finally, the latents are updated using a deterministic DDIM scheduler step [34]:
$z_{t-1}=\sqrt{\alpha_{t-1}}(\frac{z_{t}-\sqrt{1-\alpha_{t}}\cdot\hat{\epsilon}_{t}^{\text{refined}}}{\sqrt{\alpha_{t}}})+\sqrt{1-\alpha_{t-1}}\hat{\epsilon}_{t}^{\text{refined}}$ (5)
By iteratively refining the noise prediction at each timestep, this approach preserves the integrity of the content while applying the desired style transformation. This novel refinement strategy controls the incorporation of style information in the stylized image while emphasizing detailed information in the output.
4 Experiments
In this chapter, we present an experimental evaluation of our proposed method. To implement our ideas, we utilized the InstantStyle-Plus pipeline [38] as our base, built upon the Stable Diffusion XL (SDXL 1.0) model, and incorporated several modifications. For all stylization experiments, the diffusion process (inference) was performed for 10 denoising steps. The initial latent representation of the content image (or mosaic restored content image) was obtained using the inversion algorithm from [10] with 6 inversion steps. The choice of these relatively low step counts was made to evaluate the efficiency and identity preservation capabilities of our method under computationally constrained conditions, a key challenge highlighted earlier. The details of our other modifications and the experimental setup will be discussed in the following sections.
4.1 Effect of Image Captioning on Identity Preservation
During our evaluation of the InstantStyle-Plus pipeline, we observed that the base pipeline employs BLIP image captioning [24] to generate text prompts for reconstructing the structural details of the content image. Our experiments revealed that ControlNet modules, such as Canny and Tile ControlNet, can effectively preserve the content structure without relying on these descriptive prompts Figure 3. While generated prompts are intended to guide structural reconstruction, our findings show that using image captioning can sometimes lead to the generation of misleading or incorrect visual elements, due to generalized or inaccurate textual descriptions. For example, in Figure 4, the caption generated by BLIP ’a man wearing glasses is smiling in an office’ incorrectly suggests visible teeth, although the subject’s teeth are not shown in the original image. This misrepresentation causes the model to hallucinate details, such as artificially adding teeth, leading to a loss of identity fidelity. Furthermore, using image captioning models increases both memory usage and inference time, despite producing results that are visually similar to those obtained without captioning.
Figure 3: Comparison With and Without Image Captioning Prompts in the InstantStyle-Plus Pipeline. The results show that both settings produce similar outputs, with minimal visual differences.
Figure 4: The caption “a man wearing glasses is smiling in an office” generated by BLIP. The word smiling misleads the model and causes incorrect teeth generation.
Figure 5: Comparison of Style Transfer Results Using Single vs. Mosaic-Style References. Using a mosaic-style reference composed of multiple images from the same stylistic domain reduces identity and background leakage compared to a single style reference, thereby enhancing facial identity preservation and achieving a more consistent color distribution.
Figure 6: This figure showcases the outcomes of the InstantStyle-Plus pipeline with and without our proposed Mosaic Restored Content Image technique. Our approach demonstrates notable improvements in preserving individual identities, particularly in images with multiple faces or small facial regions, while maintaining a high level of stylization quality.
4.2 Effect of Mosaic Restored Content Image on Identity Preservation
The initial prompt engineering strategy, while effective for stylization, demonstrated limitations in preserving facial identity, particularly when targeting strong artistic styles. For instance, employing style-specific keywords (e.g., ”comic art,” ”graphic novel art,” ”vibrant colors,” ”highly detailed”) for comic-style generation yielded highly stylized outputs but often at the cost of significant facial feature distortion, thereby compromising subject identity.
To mitigate this critical issue, we introduced a ”Mosaic Restored Content Image” as an input to the model. As qualitatively illustrated in Figure 6 , this technique substantially improves identity preservation, especially in challenging scenarios involving multiple subjects or diminutive facial regions within the image. Compared to the baseline pipeline [38], our proposed method exhibits a marked enhancement in maintaining identity fidelity while concurrently achieving the desired level of stylization.
For a rigorous quantitative assessment of our method’s identity preservation capabilities, we conducted experiments on the IMDB dataset. Our evaluation strategy for identity preservation was informed by methodologies presented in prior work on facial attribute manipulation [28], which emphasizes the importance of maintaining subject identity alongside attribute modification. We partitioned the IMDB dataset into three categories based on the face-to-image area ratio: Category 1 (facial area ¡ 10% of total image area), Category 2 (facial area between 10-20%), and Category 3 (facial area ¿ 20%). Facial regions were automatically detected, cropped, and aligned using established facial landmark detection algorithms [35]. Identity preservation was quantified by computing the cosine similarity between the deep feature embeddings of the original content images and their stylized counterparts. These embeddings were extracted using a pre-trained WebFace recognition model [51].
Table 1: Cosine Similarity Comparison between Our Model and ISP for Different Face Sizes.
Category Our Model ISP Improvement (%)
Category 1
Anime 0.32 0.169 89.94%
Comic 0.38 0.143 165.03%
Sketch 0.581 0.201 188.06%
Category 2
Anime 0.422 0.365 15.60%
Comic 0.524 0.292 79.45%
Sketch 0.673 0.417 61.43%
Category 3
Anime 0.564 0.698 -19.21%
Comic 0.637 0.646 -1.39%
Sketch 0.727 0.704 3.27%
Table 1 presents a comparative analysis of the cosine similarity scores achieved by our proposed model versus the InstantStyle-Plus pipeline in three distinct artistic styles: anime, comic and sketch. The results compellingly demonstrate that our approach significantly outperforms the baseline in terms of identity preservation, most notably for smaller facial regions (Categories 1 and 2). This substantial enhancement is directly attributable to the efficacy of the Mosaic Restored Content Image technique, which robustly maintains salient facial characteristics even under conditions of small facial scale or significant camera-to-subject distance. For Category 3 (faces ¿20% of total image area), the results are mixed. Although our method demonstrates an improvement in the sketch style (+3.27%), the InstantStyle-Plus baseline performs better for the anime (-19. 21%) and comic (-1. 39%) styles in terms of cosine similarity. This suggests that while our technique provides substantial benefits for smaller, more challenging facial regions (Categories 1 and 2), its advantages are less pronounced for larger faces where the input facial detail is already high.
4.3 Mosaic-Style Reference for Mitigating Identity and Background Leakage
In the base pipeline, we observed that selecting a single style reference image containing a distinct identity and background often leads to unintended “leakage” of both facial characteristics and environmental details in the final stylized output. Specifically, the stylized image may inherit the identity of the person depicted in the style reference, along with its background elements and color distribution. To mitigate this issue, we propose to construct a mosaic-style reference composed of multiple images from the same stylistic domain (e.g., anime, 3D, or sketch). By diversifying the source of color distributions and structural cues, this mosaic-based approach dilutes the influence of any single image, thus reducing identity and background leakage. As illustrated in Figure 5, our method preserves the identity in the content image more effectively while maintaining a richer and more consistent color distribution in the generated output.
4.4 Impact of the Content Consistency Loss Hyperparameter on Stylization Results
Figure 7 illustrates the impact of varying the hyperparameter $\lambda_{c}$ in the content consistency loss on the preservation of fine details in the stylized output (equation  4). Our experiments reveal that increasing $\lambda_{c}$ enhances the retention of subtle content features. for example, the teeth of the girl in the image become progressively clearer as $\lambda_{c}$ is raised. This improvement in detail fidelity demonstrates the effectiveness of our noise refinement strategy in maintaining structural information. However, when $\lambda_{c}$ exceeds approximately 1200, the images exhibit noticeable color saturation, indicating an overemphasis on content preservation that compromises the intended stylistic effect. These findings underscore the critical importance of optimizing $\lambda_{c}$ to achieve a balanced trade-off between style transfer and content fidelity.
Figure 7: Effect of Varying the Content Consistency Loss on the Stylized Output, Emphasizing Attention to Fine Details.
5 Conclusion
In this work, we presented a novel, training-free framework for identity-preserving stylized image generation using diffusion models. Our work specifically addresses the challenge of maintaining identity integrity in challenging scenarios, such as images with diminutive facial regions or multiple subjects. To this end, we introduced two primary technical contributions: first, the ”Mosaic Restored Content Image” technique, which demonstrably enhances identity retention in stylized outputs; and second, a training-free content consistency loss designed to improve the preservation of fine-grained content details by focusing the model’s attention on the original image during stylization. Our findings confirm that integrating these content-aware refinements within the diffusion process leads to significant improvements in both identity preservation and overall stylistic control. We also highlighted the crucial role of text prompts in achieving desired stylization levels; specifically, strategic selection of style-related keywords was shown to effectively guide the generation towards high-quality artistic outputs. In summary, our proposed approach provides an effective and efficient solution for generating stylized images that faithfully preserve subject identity without the need for additional model training or fine-tuning. Future research could extend this work by further developing adaptive noise refinement techniques and exploring advanced prompt engineering strategies to achieve even greater control over style and content in diffusion-based image synthesis.